OpenStudy (anonymous):

Use inequality symbols to express each of the following statements: a. the number of children (variable N) is between 30 and 50 inclusively b. the sum of points (variable S) on a pair of fair die is not less than 7 c. X is greater than or equal to -4 but less than 3 d. P is at most 5

OpenStudy (dangerousjesse):

Do you know what inclusively means? :)

OpenStudy (anonymous):

It means that those numbers in this instance, are included in the total?

OpenStudy (dangerousjesse):

Yep, so can you write that out as an inequality?

OpenStudy (anonymous):

Yeah because it's in between those two numbers

OpenStudy (dangerousjesse):

I meant can you show me what you got for that :P

OpenStudy (anonymous):

oh okay I got this 30 > N > 50

OpenStudy (dangerousjesse):

\[30\le n \le 50\] :)

OpenStudy (anonymous):

oh okay oops

OpenStudy (dangerousjesse):

And two is saying that the sum is greater than or equal to seven

OpenStudy (dangerousjesse):

No problem

OpenStudy (dangerousjesse):

C is \(-4\le x < 3\)

OpenStudy (dangerousjesse):

And d is saying that P is less than or equal to 5

OpenStudy (anonymous):

okay awesome I got b and c

OpenStudy (anonymous):

and D. P < or equal to 5